OpenStudy (anonymous):

Find the non-extraneous solutions of the square root of x+6-5=x+1

OpenStudy (ikram002p):

non-extraneous ?

OpenStudy (anonymous):

yeah thats what the q says

OpenStudy (anonymous):

here are the answer choices: x = −2 x = −6 x = −6 and x = −2 x = 2

OpenStudy (ikram002p):

is it \(\sqrt{ ( x+6)}-5=x+1\) ?

OpenStudy (ikram002p):

\(\sqrt{ ( x+6)} =x+6\) so x=-6, -4 not sure what is non-extraneous @ganeshie8 ?

OpenStudy (ikram002p):

sorry i ment =-6,-2

OpenStudy (anonymous):

yeah but without the peranthasis